# Some new cases of Zilber–Pink in $Y(1)^{3}$

### v1

@@ 4.5.1. Reductions
After possibly relabeling the coordinates, we can assume that $\overline{C}\setminus C$ contains a point $s_{0}:=(b_{1},b_{2},\infty)$ with $(b_{1},b_{2})\in Y(1)^{2}$ modular. After possibly removing finitely many points from $C$, namely, those belonging to the special hyperplanes defined by setting a coordinate to 0 or $1728$, we can assume that $C$ is contained in $S^{3}_{K}$. After possibly replacing $K$ with a finite extension and $\overline{C}$ with a finite étale cover, we obtain a covering datum
$\mathcal{D}:=(\mathfrak{G}^{\prime}_{1},\mathfrak{G}^{\prime}_{2},\mathfrak{G}^{\prime}_{3},\tilde{\mathfrak{C}}^{\prime\prime},\tilde{\mathfrak{C}}^{\prime},\tilde{C}^{\prime},\tilde{C},\nu,x,s_{1},\dots,s_{\ell},\sigma_{1},\dots,\sigma_{\ell})$
@@
4.5.2. Setup
For $I:=\{1,2\}$, let $\Lambda\subset\{1,\ldots,\ell\}\times I$ be a set of pullback representatives for $(\mathcal{D},I)$. As in [DO22, Rem. 5.3], we have $j,1]\neq[j,2]$ for all $j\in\{1,\ldots,\ell\}$. Let $\tilde{\mathfrak{C}}\subset\tilde{\mathfrak{C}}$$^{\prime\prime}$ denote the maximal Zariski open subset such that $\mathfrak{G}_{\lambda}:=\mathfrak{G}^{\prime}_{\lambda}\times_{\tilde{\mathfrak{C}}^{\prime\prime}}\tilde{\mathfrak{C}}$ is an elliptic scheme for all $\lambda\in\Lambda$ and replace $\tilde{C}$ with $\tilde{\mathfrak{C}}_{K}$ (note that we still have $s_{j}\in\tilde{C}(K)$ for all $j\in\{1,\ldots,\ell\})$. For $\lambda\in\Lambda$, define $\mathcal{E}_{\lambda}:=\mathcal{E}^{\prime}_{\lambda}|_{\tilde{C}}$.
To simplify notation, remove all $\tilde{\cdot}$ decorations. By [DO22, Lemma 5.4], for any $j\in\{1,\ldots,\ell\}$, there exists an isogeny
$\mathcal{E}_{[j,1],s_{1}}=\mathcal{E}^{\prime}_{[j,1],s_{1}}\to(\sigma^{*}_{j}\mathcal{E}^{\prime}_{1})_{s_{1}}=\mathcal{E}^{\prime}_{1,s_{j}}\to\mathcal{E}^{\prime}_{2,s_{j}}=(\sigma^{*}_{j}\mathcal{E}^{\prime}_{2})_{s_{1}}\to\mathcal{E}^{\prime}_{[j,2],s_{1}}=\mathcal{E}_{[j,2],s_{1}}$
of elliptic curves, which we denote $f_{j}$. We choose the $f_{j}$ compatibly, so that, if $j,i]=[j^{\prime},i]$, then the induced diagram of isogenies commutes. Furthermore, after possibly replacing $K$ with a finite extension, we may assume that the $f_{j}$ are defined over $K$ and, after possibly removing finitely many points from $C$, that $H^{1}_{DR}(\mathcal{E}_{[j,i]}/C)$ is free. For each $\lambda=[j,i]\in\Lambda$, choose bases $\{\omega_{\lambda},\eta_{\lambda}\}$ for $H^{1}_{DR}(\mathcal{E}_{\lambda}/C)$ as in Section 3, with $s_{1}\in C(K)$ playing the role of $s_{0}$ and $f_{j}$ the role of $f_{0}$. (This is possible due to the compatibility of the $f_{j}$.)
Let $F_{\lambda mn}\in K[[X]]$ ($1\leq m,n\leq 2$) denote the G-functions associated with $\mathcal{E}_{\lambda}\to C$, $x$ and $\{\omega_{\lambda},\eta_{\lambda}\}$ (see Section 2.6). We define
$\mathcal{G}=\{F_{\lambda mn}:\lambda\in\Lambda,\ 1\leq m,n\leq 2\},$
and, for each place $v$ of $K$, we define
$R_{v}=\min\{1,R(F^{v\textrm{-}\mathrm{an}}):F\in\mathcal{G}\}.$
Let $((r_{v},U_{v,1},\dotsc,U_{v,\ell})_{v},H)$ denote a neighbourhood system for $(C^{\prime},x)$ and define $\mathcal{G}_{H}:=\mathcal{G}\cup\{H\}$. For each $v\in\Sigma_{K}$, define
$R^{\leq}_{v}:=\min\{R_{v},R(H^{v\textrm{-}\mathrm{an}})\}\leq r_{v},$
and, for each $j\in\{1,\ldots,\ell\}$, let $U^{\prime}_{v,j}=(x^{v\textrm{-}\mathrm{an}}|_{U_{v,j}})^{-1}(D(0,R^{\leq}_{v},K_{v}))$. By Lemma 9, after possibly modifying $H$ (in other words, altering finitely many of the $r_{v}$), we can ensure that, for any $j\in\{1,\ldots,\ell\}$, the data of
$C^{\prime},\mathfrak{C}^{\prime\prime},\mathfrak{G}^{\prime}_{\lambda},C,\mathcal{E}_{\lambda},\mathfrak{C},\mathfrak{G}_{\lambda},s_{1},f_{j},\{\omega_{\lambda},\eta_{\lambda}\},x,F_{\lambda mn},\Delta_{v}:=U^{\prime}_{v,1},$
for $\lambda=[j,1],[j,2]$ and $v\in\Sigma_{K}$, satisfy the assumptions of Section 3.
@@ 4.5.3. Constructing relations
Consider $s\in C(\overline{\mathbb{Q}})$ as above and replace $K$ with $K(s)$. For any $j\in\{1,\ldots,\ell\}$, we obtain an isogeny
$\mathcal{E}_{[j,1],\sigma_{j}^{-1}(s)}\to(\sigma^{*}_{j}\mathcal{E}_{1})_{\sigma^{-1}_{j}(s)}=\mathcal{E}_{1,s}\to\mathcal{E}_{2,s}=(\sigma^{*}_{j}\mathcal{E}_{2})_{\sigma^{-1}_{j}(s)}\to\mathcal{E}_{[j,2],\sigma_{j}^{-1}(s)},$
the outer isogenies being afforded to us, again, by [DO22, Lemma 5.4]. Moreover, for any $v\in\Sigma_{K}$ such that $|x(s)|_{v}<R^{\leq}_{v}$, we have $s^{v\textrm{-}\mathrm{an}}\in U^{\prime}_{v,j}$ for some $j\in\{1,\ldots,\ell\}$. As in [DO22, Rem. 5.6], we deduce that $\sigma^{-1}_{j}(s)^{{v\textrm{-}\mathrm{an}}}\in U^{\prime}_{v,1}=\Delta_{v}$. Hence, by Proposition 11, there exists a $v$-adic relation $Q_{v}$ of degree at most $2$ between the evaluations at $x(\sigma^{-1}_{j}(s))=x(s)$ of the $F_{[i,j]mn}$ for $i,m,n\in\{1,2\}$ which is not contained in the ideal $I$ of $\overline{\mathbb{Q}}[X_{imn}:1\leq i,m,n\leq 2]$ generated by the elements
@@
Let $\Sigma_{s}$ denote the set of $v\in\Sigma_{K}$ satisfying $|x(s)|_{v}<R^{\leq}_{v}$ and let
@@
Then $Q$ is a global relation between the evaluations at $x(s)$ of the elements of $\mathcal{G}$ (and, hence, $\mathcal{G}_{H}$). To constrain its degree, we observe that $v\in\Sigma_{s}$ implies that $v\in\Sigma_{K,\infty}$ or $v\in\Sigma_{K,f}$ and, writing $\mathfrak{s}_{1}$, $\mathfrak{s}_{1}(j)$, and $\mathfrak{s}(j)$ for the canonical extensions of $s_{1}$, $\sigma^{-1}_{j}(s_{1})$, and $\sigma^{-1}_{j}(s_{1})$, respectively, we have an isogeny
$\mathfrak{G}^{\prime}_{1,\mathfrak{s}_{1},k_{v}}=\mathfrak{G}^{\prime}_{[j,1],\mathfrak{s}_{1}(j),k_{v}}\cong\mathfrak{G}^{\prime}_{[j,1],\mathfrak{s}(j),k_{v}}\to\mathfrak{G}^{\prime}_{[j,3],\mathfrak{s}(j),k_{v}}\cong\mathfrak{G}^{\prime}_{[j,3],\mathfrak{s}_{1}(j),k_{v}}=\mathfrak{G}^{\prime}_{3,\mathfrak{s}_{1},k_{v}}=\mathbb{G}_{m,k_{v}},$
where we use the fact that $\sigma^{-1}_{j}(s)^{{v\textrm{-}\mathrm{an}}}\in\Delta_{v}$. In other words, $v\in\Sigma_{s}$ implies that $v\in\Sigma_{K,\infty}$ or $v\in\Sigma_{K,f}$ and $v$ is a place of bad reduction for $\mathcal{E}_{1,s_{0}}$.
@@ 4.5.4. Nontriviality
By [DO22, Lemma 2.9], it suffices to show that $Q$ is a non-trivial relation between the elements of $\mathcal{G}$ (as opposed to $\mathcal{G}_{H}$). Therefore, let $J$ denote the (homogeneous) ideal of
@@
comprising the functional relations between the elements of $\mathcal{G}$. We need to show that $Q\notin J$. By the following lemma, it suffices to show that $Q_{v}\notin J$ for all $v\in\Sigma_{s}$.
Lemma 15.
@@
To ease notation, we set $k:=|\Lambda|$. Choose $v\in\Sigma_{K,\infty}$. The $F_{\lambda mn}$ give rise to a function $D(0,R^{\leq}_{v},\mathbb{C})\to\mathbf{SL}^{k}_{2}(\mathbb{C})$, and we let $\Gamma$ denote the graph of this function.
Lemma 16.
The set $\Gamma$ is $\mathbb{C}$-Zariski dense in $\mathbb{A}^{1}\times\mathbf{SL}_{2}^{k}$.
@@
It suffices to show that
$\operatorname{trdeg}_{\mathbb{C}(X)}\mathbb{C}(X)(F_{\lambda mn}:\lambda\in\Lambda,\ 1\leq m,n\leq 2)\geq 3k.$
To that end, denote by $\Omega_{v}:\Delta_{v}\to\mathrm{M}_{2g}(\mathbb{C})$ the function giving rise to the $F_{\lambda mn}$, as in Section 2.4, and we write $\Omega_{v\lambda mn}$ for its components. Then the previous inequality is clearly implied by
$\operatorname{trdeg}_{\mathbb{C}(C)}\mathbb{C}(C)(\Omega_{v\lambda mn}:\lambda\in\Lambda,\ 1\leq m,n\leq 2)\geq 3k.$
As in Remark 7, if we define $\Omega_{v}$ with suitable horizontal bases for $H^{1}_{DR}(\mathcal{E}_{\lambda}/C)^{v\textrm{-}\mathrm{an}}(\Delta_{v})$, the $\Omega_{v\lambda mn}$ are the (non-zero) entries of a period matrix for $\mathcal{E}_{\lambda}\to C$. Therefore, the geometric André–Grothendieck period conjecture, proved in this case by Ayoub [Ayo15] and Nori (unpublished)—see [BT25, Thm. 1.1] and the remarks thereafter—yields exactly this inequality. ∎
By definition, we have $\Gamma\subset V(J)$. Thus, since $J$ is homogeneous, Lemma 16 implies that
$V(J)=\mathbb{A}^{1}\times\{(g_{1},\ldots,g_{k})\in\mathbf{GL}_{2}^{k}(\mathbb{C}):\det(g_{1})=\ldots=\det(g_{k})\}.$
Since $J$ is prime and therefore radical, we deduce that $J$ is the ideal generated by the elements
@@
for $\lambda_{1},\lambda_{2}\in\Lambda$.
Now let $v\in\Sigma_{s}$. The fact that $Q_{v}\notin J$ now follows from the fact that $Q_{v}$ is not contained in the ideal $I$ of Proposition 11 (by faithful flatness of a polynomial ring over its ring of coefficients, for example). ∎
4.6. Proof of Theorem 6
Consider the situation in Theorem 6. In particular, let $s_{0}:=(b_{1},b_{2},b_{3})\in C$ be the modular point referred to in the statement, and let $K$ denote a number field over which $C$ and $s_{0}$ are defined. Let $\epsilon>0$.

### v2

@@ 4.5.1. Reductions
After possibly relabeling the coordinates, we can assume that $C\setminus C_{0}$ contains a point $s_{0}:=(b_{1},b_{2},\infty)$ with $\Phi_{N}(b_{1},b_{2})=0$ for some $N\in\mathbb{N}$. After possibly removing finitely many points from $C_{0}$, namely, those belonging to the special hyperplanes defined by setting a coordinate to 0 or $1728$, we can assume that $C_{0}$ is contained in $S^{3}_{K}$. Define $\mathcal{E}_{i,0}$ and $\mathcal{E}_{i}$ as in Sections 4.1 and 4.2, respectively.
Thus, after possibly replacing $K$ with a finite extension and $C$ with a finite étale cover, we obtain a covering datum
$\mathcal{D}:=(\mathfrak{G}_{1},\mathfrak{G}_{2},\mathfrak{G}_{3},{\mathfrak{C}}^{\prime},{\mathfrak{C}},\tilde{C},\tilde{C}_{0},\nu,x,s_{1},\dots,s_{\ell},\sigma_{1},\dots,\sigma_{\ell})$
for $(C,s_{0})$ (as defined in Section 4.2). Fix a Weil height $h$ on $\tilde{C}$. By standard properties of heights, it suffices for us to show that there exist constants $c_{7}$ and $c_{8}$} such that, for any modular point $s:=(x_{1},x_{2},x_{3})\in C_{0}$, and any $\tilde{s}\in\nu^{-1}(s)\in\tilde{C}_{0}$, we have
@@
4.5.2. Pullback representatives
For $I:=\{1,2\}$, let $\Lambda\subset\{1,\ldots,\ell\}\times I$ be a set of pullback representatives for $(\mathcal{D},I)$. As in [DO25, Rem. 5.3], we have $k,1]\neq[k,2]$ for all $k\in\{1,\ldots,\ell\}$. Define $\mathcal{E}_{\lambda}$ and $\mathfrak{G}_{\lambda}$ as in Section 4.3 and let ${\mathfrak{C}}_{0}\subset{\mathfrak{C}}^{\prime}$ denote a Zariski open subset such that $s_{1}\in\mathfrak{C}_{0}(K)$ and $\mathfrak{G}_{\lambda}\times_{{\mathfrak{C}}^{\prime}}{\mathfrak{C}}_{0}$ is an elliptic scheme for all $\lambda\in\Lambda$. Relabel $C:=\tilde{C}$ and $C_{0}:=\tilde{C}_{0}$, and replace ${C}_{0}$ with ${\mathfrak{C}}_{0,K}$. Write $\mathcal{E}_{\lambda,0}:=\mathcal{E}_{\lambda}\times_{C}C_{0}$. After possibly replacing $K$ with a finite extension and possibly removing finitely many points from $C_{0}$, we may assume that $H^{1}_{DR}(\mathcal{E}_{\lambda,0}/C_{0})$ is free for all $\lambda\in\Lambda$.
4.5.3. Isogeny data
By [DO25, Lemma 5.4] and the hypothesis that $(b_{1},b_{2})$ is modular, for each $k\in\{1,\ldots,\ell\}$, there exist isogenies
$\mathcal{E}_{[k,1],s_{1}}\to(\sigma^{*}_{k}\mathcal{E}_{1})_{s_{1}}=\mathcal{E}_{1,s_{k}}=\mathcal{E}_{b_{1}}\to\mathcal{E}_{b_{2}}=\mathcal{E}_{2,s_{k}}=(\sigma^{*}_{k}\mathcal{E}_{2})_{s_{1}}\to\mathcal{E}_{[k,2],s_{1}}$
of elliptic curves over $\overline{K}$. We want to choose isogenies $\mathcal{E}_{[k,1],s_{1}}\to\mathcal{E}_{[k,2],s_{1}}$, but we shall require certain compatibility conditions between these isogenies for different $k$. We achieve this using the following construction.
Consider the undirected graph $G$ with vertex set $V(G)=\Lambda$ and edge set $E(G)$ comprising pairs of vertices of the form $[k,1]$ and $[k,2]$ for some $k\in\{1,\ldots,\ell\}$. Note that, for each $e\in E(G)$, the index $k\in\{1,\ldots,\ell\}$ giving rise to that edge is not necessarily unique. Thus, for each $e\in E(G)$, we assign such a $k:=k(e)$, and we define $e(i):=[k(e),i]$ (which is an endpoint of $e$).
Choose an isogeny $f_{e}\colon\mathcal{E}_{e(1),s_{1}}\to\mathcal{E}_{e(2),s_{1}}$ for each edge $e$ as follows.
  1. (1)
    Let $T$ denote a spanning forest for $G$. That is, for each connected component of $G$, choose a spanning tree, and let $T$ be the union of these spanning trees.
  2. (2)
    For each $e\in E(T)$, choose $f_{e}:\mathcal{E}_{e(1),s_{1}}\to\mathcal{E}_{e(2),s_{1}}$ arbitrarily.
  3. (3)
    For each remaining $e\in E(G)$, there exists a unique sequence of edges $e_{1},\ldots,e_{r}\in E(T)$ connecting $e(1)$ to $e(2)$; to each edge $e_{i}$ define
@@ 4.5.3. Isogeny data
    and define $f_{e}:\mathcal{E}_{e(1),s_{1}}\to\mathcal{E}_{e(2),s_{1}}$ to be the composition $g_{e_{r}}\circ\cdots\circ g_{e_{1}}$.
After possibly replacing $K$ with a finite extension, we may assume that the $f_{e}$ are defined over $K$. For each $\lambda\in\Lambda$, we choose bases $\{\omega_{\lambda},\eta_{\lambda}\}$ for $H^{1}_{DR}(\mathcal{E}_{\lambda,0}/C_{0})$ such that, for any $e\in E(G)$, we have
(13) $\omega_{e(1),s_{1}}=\mu_{e}f_{e}^{*}\omega_{e(2),s_{1}}\text{ and }\eta_{e(1),s_{1}}=\mu_{e}f_{e}^{*}\omega_{e(2),s_{1}},$
for some $\mu_{e}\in\mathbb{Q}$ (this is the analogue of condition (3) of Section 3, with $s_{1}\in C(K)$ playing the role of $s_{0}$ and $f_{e}$ the role of $f_{0}$). In order to do this, consider a connected component $G^{\prime}\subset G$. Choose a vertex $\lambda_{0}\in V(G^{\prime})$ and choose an arbitrary basis $\{\omega_{\lambda_{0}},\eta_{\lambda_{0}}\}$ for $H^{1}_{DR}(\mathcal{E}_{\lambda_{0},0}/C_{0})$. Working outwards from $\lambda_{0}$, following the edges of the tree $G^{\prime}\cap T$, we then choose a basis for $H^{1}_{DR}(\mathcal{E}_{\lambda,0}/C_{0})$ for each $\lambda\in V(G^{\prime})$, satisfying (13) for all $e\in E(G^{\prime}\cap T)$ (for $e\in E(G^{\prime}\cap T)$, we may choose $\mu_{e}$ arbitarily; for example, we may choose $\mu_{e}=1$ for all $e\in E(G^{\prime}\cap T)$). Finally, observe that the chosen bases satisfy (13) for all $e\in E(G^{\prime})\setminus E(T)$, due to the construction of the $f_{e}$ above. (Note that we are no longer free to choose the values of $\mu_{e}$ for $e\in E(G^{\prime})\setminus E(T)$, and we might not have $\mu_{e}=1$ for all such $e$ because any edge $e_{i}$ which was traversed in the reverse direction while constructing $f_{e}$ will introduce a factor of $\deg(f_{e_{i}})$ in $\mu_{e}$.)
Let $F_{\lambda mn}\in K[\![X]\!]$ ($1\leq m,n\leq 2$) denote the G-functions associated with $\mathcal{E}_{\lambda,0}\to C_{0}$, $x$, $s_{1}$ and $\{\omega_{\lambda},\eta_{\lambda}\}$ (see Section 2.6) and define
$\mathcal{G}=\{F_{\lambda mn}:\lambda\in\Lambda,\ 1\leq m,n\leq 2\}.$
Let $\mathcal{S}:=((r_{v},U_{v,1},\dotsc,U_{v,\ell})_{v})$ denote a neighbourhood system for $(C,x)$ and let $H$ denote an auxillary G-function for $\mathcal{S}$. Define $\mathcal{G}_{H}:=\mathcal{G}\cup\{H\}$. For each $v\in\Sigma_{K}$, define
$R_{v}:=\min\{1,R(H^{v\textrm{-}\mathrm{an}}),R(F^{v\textrm{-}\mathrm{an}}):F\in\mathcal{G}\}\leq r_{v}.$
For each $k\in\{1,\ldots,\ell\}$, let $U_{R_{v},k}=(x^{v\textrm{-}\mathrm{an}}|_{U_{v,k}})^{-1}(D(0,R_{v},K_{v}))$, and let $\Delta_{v}:=U_{R_{v},1}$. Then, for each $e\in E(G)$,
$\{C,\mathfrak{C},\mathfrak{G}_{e(i)},C_{0},\mathcal{E}_{e(i),0},\mathfrak{C}_{0},s_{1},f_{e},\{\omega_{e(i)},\eta_{e(i)}\},x,F_{e(i)mn},R_{v},\Delta_{v}:1\leq i,m,n\leq 2,v\in\Sigma_{K}\}$
is an isogeny datum (see Section 3).
@@ 4.5.4. Constructing relations
Consider $s\in C_{0}$ with $\nu(s)$ modular and replace $K$ with $K(s)$. There exists an isogeny $\mathcal{E}_{1,s}\to\mathcal{E}_{2,s}$. (By the definition of modular points, there also exists an isogeny $\mathcal{E}_{1,s}\to\mathcal{E}_{3,s}$, but the latter isogeny is not required for this step of the proof; it will be used in Section 4.5.5 below.) Hence, for each $k\in\{1,\ldots,\ell\}$, there is an isogeny
$\mathcal{E}_{[k,1],\sigma_{k}^{-1}(s)}\to(\sigma^{*}_{k}\mathcal{E}_{1})_{\sigma^{-1}_{k}(s)}=\mathcal{E}_{1,s}\to\mathcal{E}_{2,s}=(\sigma^{*}_{k}\mathcal{E}_{2})_{\sigma^{-1}_{k}(s)}\to\mathcal{E}_{[k,2],\sigma_{k}^{-1}(s)},$
the outer isogenies being afforded to us, again, by [DO25, Lemma 5.4]. After replacing $K$ with an extension of bounded degree, we can and do assume these are $K$-isogenies.
Now, for any $v\in\Sigma_{K}$ such that $|x(s)|_{v}<R_{v}$, we have $s^{v\textrm{-}\mathrm{an}}\in U_{R_{v},k}$ for some $k\in\{1,\ldots,\ell\}$ (which may depend on $v$, of course). As in [DO25, Rem. 5.6], we deduce that $\sigma^{-1}_{k}(s)^{{v\textrm{-}\mathrm{an}}}\in U_{R_{v},1}=\Delta_{v}$. Let $e\in E(G)$ be such that $e(i)=[k,i]$ for $i\in\{1,2\}$ (such an $e$ exists by the definition of $G$). Hence, by Proposition 3.2 applied to
$\{C,\mathfrak{C},\mathfrak{G}_{e(i)},C_{0},\mathcal{E}_{e(i),0},\mathfrak{C}_{0},s_{1},f_{e},\{\omega_{e(i)},\eta_{e(i)}\},x,F_{e(i)mn},R_{v},\Delta_{v}:1\leq i,m,n\leq 2,v\in\Sigma_{K}\},$
there exists a $v$-adic relation $Q_{v}$ of degree at most $2$ between the evaluations at $x(\sigma^{-1}_{k}(s))=x(s)$ of the $F_{e(i)mn}$ for $i,m,n\in\{1,2\}$, with the property that $Q_{v}$ is not contained in the ideal $I$ of $\overline{\mathbb{Q}}[X_{imn}:1\leq i,m,n\leq 2]$ generated by the elements
@@
Let $\Sigma_{s}$ denote the set of $v\in\Sigma_{K}$ satisfying $|x(s)|_{v}<R_{v}$ and let
@@
Then $Q$ is a global relation between the evaluations at $x(s)$ of the elements of $\mathcal{G}_{H}$, thanks to the definition of $R_{v}$.
4.5.5. Constraining the degree of $Q$
Consider a place $v\in\Sigma_{s}\cap\Sigma_{K,f}$ and the corresponding index $k$ such that $s^{{v\textrm{-}\mathrm{an}}}\in U_{R_{v},k}$. Write $\mathfrak{s}_{k}$ and $\mathfrak{s}$ for the $\mathcal{O}_{K}$-extensions of $s_{k}$ and $s$, respectively, in $\mathfrak{C}^{\prime}$. Since $s$ is modular, there exists an isogeny $\mathcal{E}_{1,s}\to\mathcal{E}_{3,s}$ and hence an isogeny between the Néron models $\mathfrak{G}_{1,\mathfrak{s}}\to\mathfrak{G}_{3,\mathfrak{s}}$. Since $s^{{v\textrm{-}\mathrm{an}}}\in U_{R_{v},k}$ and $U_{R_{v},k}$ is a simple neighbourhood, we have $\mathfrak{s}(\mathfrak{p}_{v})=\mathfrak{s}_{k}(\mathfrak{p}_{v})$. Therefore, we have an isogeny
$\mathfrak{G}_{1,\mathfrak{s}_{k},k_{v}}\cong\mathfrak{G}_{1,\mathfrak{s},k_{v}}\to\mathfrak{G}_{3,\mathfrak{s},k_{v}}\cong\mathfrak{G}_{3,\mathfrak{s}_{k},k_{v}}=\mathbb{G}_{m,k_{v}}.$
Hence, $\mathcal{E}_{b_{1}}=\mathcal{E}_{1,s_{k}}$ has bad reduction at $v$.
@@ 4.5.5. Constraining the degree of $Q$
We conclude that Theorem 1.7 follows from [And89, VII, Thm. 5.2] (cf. [DO25, Thm. 2.8]) if we can show that $Q$ is nontrivial. This will be the subject of the next section.
@@ 4.5.6. Nontriviality
By [DO25, Lemma 2.9], it suffices to show that $Q$ is a non-trivial relation between the elements of $\mathcal{G}$ (as opposed to $\mathcal{G}_{H}$). Therefore, let $J$ denote the (homogeneous in $X_{\lambda mn}$) ideal of
@@
comprising the functional relations between the elements of $\mathcal{G}$. We need to show that $Q$ does not belong to the specialisation of $J$ at $x(s)$.
Lemma 4.1.
@@
To ease notation, we set $r:=|\Lambda|$. Choose $v\in\Sigma_{K,\infty}$. The $F_{\lambda mn}$ give rise to a holomorphic function $D(0,R_{v},\mathbb{C})\to\mathbf{SL}^{r}_{2}(\mathbb{C})$, and we let $\Gamma$ denote the graph of this function.
Lemma 4.2.
The set $\Gamma$ is $\mathbb{C}$-Zariski dense in $\mathbb{A}^{1}\times\mathbf{SL}_{2}^{r}$.
@@
Denote by $\Omega_{v}:\Delta_{v}\to\mathrm{M}_{2g}(\mathbb{C})$ the function giving rise to the $F_{\lambda mn}$, as in Section 2.6, for some choice of horizontal basis of $H^{1}_{DR}(\mathcal{E}_{\lambda}/C)^{v\textrm{-}\mathrm{an}}(\Delta_{v})$. Write $\Omega_{v\lambda mn}$ for its components.
By the Legendre period relation, $\det\Omega_{v}$ is constant. It follows that $\Gamma$ is indeed contained in $\mathbb{A}^{1}\times\mathbf{SL}_{2}^{r}$.
Therefore, it suffices to show that
$\operatorname{trdeg}_{\mathbb{C}(X)}\mathbb{C}(X)(F_{\lambda mn}:\lambda\in\Lambda,\ 1\leq m,n\leq 2)\geq\dim(\mathbf{SL}_{2}^{r})=3r.$
This is clearly implied by
$\operatorname{trdeg}_{\mathbb{C}(C)}\mathbb{C}(C)(\Omega_{v\lambda mn}:\lambda\in\Lambda,\ 1\leq m,n\leq 2)\geq 3r.$
As in Remark 2.2, if we choose suitable horizontal bases for $H^{1}_{DR}(\mathcal{E}_{\lambda}/C)^{v\textrm{-}\mathrm{an}}(\Delta_{v})$, then the $\Omega_{v\lambda mn}$ are the entries of a period matrix for $\mathcal{E}_{\lambda}\to C$. Therefore, the geometric André–Grothendieck period conjecture, proved in this case by Ayoub [Ayo15] and Nori (unpublished)—see [BT25, Thm. 1.1] and the remarks thereafter—yields exactly this inequality. Indeed, by [BT25, Cor. 4.17] the relative motivic Galois group $\mathbf{G}_{\rm mot}(M/\mathbb{C})$ is always equal to the Zariski closure of the monodromy of the Betti local system of $M$. In our case, the latter is $\mathbf{SL}_{2}^{r}$ because the $\mathcal{E}_{\lambda}$ are not generically CM or generically isogenous, so $\dim\mathbf{G}_{\rm mot}(M/\mathbb{C})=3r$. ∎
By definition, $V(J)$ is the cone over the Zariski closure of $\Gamma$. Thus, Lemma 4.2 implies that
$V(J)=\mathbb{A}^{1}\times\{(g_{1},\ldots,g_{r})\in M_{2}(\mathbb{C})^{r}:\det(g_{1})=\ldots=\det(g_{r})\}.$
Since $J$ is prime and therefore radical, we deduce that $J$ is the ideal of
$\overline{\mathbb{Q}}[X][X_{\lambda mn}:\lambda\in\Lambda,\ 1\leq m,n\leq 2]$
generated by the elements
@@
for $\lambda_{1},\lambda_{2}\in\Lambda$. The specialisation $J_{0}$ of $J$ at $x(s)$ is therefore the ideal of
$\overline{\mathbb{Q}}[X_{\lambda mn}:\lambda\in\Lambda,\ 1\leq m,n\leq 2]$
generated by these elements.
By faithful flatness of a polynomial ring over its ring of coefficients, we conclude that $J_{0}=J\cap\overline{\mathbb{Q}}[X_{\lambda mn}:\lambda\in\Lambda,\ 1\leq m,n\leq 2]$, hence $J_{0}$ is also prime. Therefore, it suffices to show that $Q_{v}\notin J_{0}$ for all $v\in\Sigma_{s}$. This follows from the fact that each $Q_{v}$ is not contained in the ideal $I$ of Proposition 3.2 (again by faithful flatness of a polynomial ring over its ring of coefficients). ∎
4.6. Proof of Theorem 1.8
Consider the situation in Theorem 1.8. In particular, let $s_{0}:=(b_{1},b_{2},b_{3})\in C$ be the modular point referred to in the statement, and let $K$ denote a number field over which $C$ and $s_{0}$ are defined. Again, to match the notation of the preceding sections, we relabel $C_{0}:=C$ and $C:=\overline{C}$.